# Quantum Dynamic Programming

### v1

@@ V.3 QDP implementation for double-bracket iterations
Now we analyze the QDP implementation using states $\rho_{n}$ in Eq. (158) as instructions for the recursion $e^{s_{n}[\hat{D},\hat{H}_{n}]}$. First, we establish the locally accurate implementation of a single recursion step, whose memory-call is of a linear Hermitian-preserving map $\mathcal{N}(\rho)=-is[\hat{D},\rho]$. Observe that
$e^{s_{n}[\hat{D},\hat{H}_{n}]}=e^{i\operatorname{Tr}[\hat{H}_{0}+\lambda\mathmybb{1}]\hat{\mathcal{N}}(\rho_{n})}\ .$ (159)
Henceforth, we rescale the flow duration $s_{n}\mapsto\operatorname{Tr}[\hat{H}_{0}+\lambda\mathmybb{1}]s_{n}$ and define our recursion as $e^{i\mathcal{N}(\rho_{n})}$.
The memory-call is then implemented by memory-usage queries to Hermitian-preserving matrix exponentiation channel ^$\hat{\mathcal{E}}^{(\mathcal{N},\rho_{n})}_{r}$, given as
^$\hat{\mathcal{E}}^{(\mathcal{N},\rho_{n})}_{r}(\sigma)=\operatorname{Tr}_{1}\left[e^{-ir\hat{N}}(\rho\otimes\sigma)e^{ir\hat{N}}\right]\ .$ (160)
@@
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{(\mathcal{N},\rho_{n})}_{1/M}\right)^{M}-\hat{\mathbf{E}}^{(\mathcal{N},\rho_{n})}\right\rVert_{\operatorname{Tr}}=\mathcal{O}\left(\frac{\lVert\hat{N}\rVert_{\infty}^{2}}{M}\right)\ ,$ (161)
from Eq. (38). The flow duration is chosen to be $s_{n}=\mathcal{O}(\lVert\hat{D}\rVert^{-1})$ as in Lemma 22, which makes the error scaling independent of the rescaling of $\hat{D}$. Therefore, one can make the locally accurate implementation, by increasing the number of copies $M$.
@@
V.4 Oblivious Schmidt decomposition
@@
$|{\psi}\rangle_{AB}=\sum_{j}\sqrt{\lambda_{j}}|{\phi_{j}}\rangle_{A}|{\chi_{j}}\rangle_{B}\ ,$ (162)
@@
$e^{i\mathcal{N}(\psi_{n})}=e^{s_{n}[\hat{D},\rho_{n}^{(A)}]}\otimes\mathmybb{1}_{B}\ ,$ (163)
@@
^$\hat{\mathcal{E}}^{(\mathcal{N},\psi_{A_{1}B_{1}})}_{r}(\sigma_{A_{2}B_{2}})=\operatorname{Tr}_{A_{1}B_{1}}\left[e^{-ir\hat{N}}(\psi_{A_{1}B_{1}}\otimes\sigma_{A_{2}B_{2}})e^{ir\hat{N}}\right]\ ,$ (164)
where $\hat{N}=is_{n}\sum_{jk}(\mu_{k}-\mu_{j})|{kj}\rangle\!\langle{jk}|_{A_{1}A_{2}}\otimes\mathmybb{1}_{B_{1}B_{2}}$ and $\{\mu_{j}\}_{j}$ are the diagonal entries of $\hat{D}$. Each query ^$\hat{\mathcal{E}}^{(\mathcal{N},\psi_{A_{1}B_{1}})}_{r}(\sigma_{A_{2}B_{2}})$ consumes a copy of $\psi_{A_{1}B_{1}}$. Furthermore, following the same calculation from the previous subsection, we obtain
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{(\mathcal{N},\psi_{A_{1}B_{1}})}_{1/M}\right)^{M}-e^{i\mathcal{N}(\psi_{n})}\right\rVert_{\operatorname{Tr}}=\mathcal{O}\left(\frac{\lVert\hat{N}\rVert_{\infty}^{2}}{M}\right)\ ,$ (165)
@@
If recursions are implemented with sufficiently good precision, Thm. 1 holds and we approach the fixed-point of the recursion, which is a diagonal matrix $\rho_{\infty}^{(A)}$ isospectral to the initial reduced state $\rho_{0}^{(A)}$. In other words, $\rho_{N}^{(A)}$ with some large number $N$ becomes close to the desired diagonal matrix
$\rho_{N}\simeq\sum_{j}\lambda_{j}|{j}\rangle\!\langle{j}|_{A}\ ,$ (166)

### v2

@@ V.3 QDP implementation for double-bracket iterations
Now we analyze the QDP implementation using states $\rho_{n}$ in Eq. (S146) as instructions for the recursion $e^{s_{n}[\hat{D},\hat{P}_{n}]}$. First, we establish the locally accurate implementation of a single recursion step, whose memory-call is of a linear Hermitian-preserving map $\mathcal{N}(\rho)=-is[\hat{D},\rho]$. Observe that
$e^{s_{n}[\hat{D},\hat{P}_{n}]}=e^{i\operatorname{Tr}[\hat{P}_{0}+\lambda\mathmybb{1}]\hat{\mathcal{N}}(\rho_{n})}\ .$ (S147)
Henceforth, we rescale the flow duration $s_{n}\mapsto\operatorname{Tr}[\hat{P}_{0}+\lambda\mathmybb{1}]s_{n}$ and define our recursion as $e^{i\mathcal{N}(\rho_{n})}$.
The memory-call is then implemented by memory-usage queries to Hermitian-preserving matrix exponentiation channel ^$\hat{\mathcal{E}}^{[\mathcal{N},\rho_{n}]}_{r}$, given as
^$\hat{\mathcal{E}}^{[\mathcal{N},\rho_{n}]}_{r}(\sigma)=\operatorname{Tr}_{1}\left[e^{-ir\hat{N}}(\rho\otimes\sigma)e^{ir\hat{N}}\right]\ .$ (S148)
@@
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{[\mathcal{N},\rho_{n}]}_{1/M}\right)^{M}-\hat{\mathbf{E}}^{[\mathcal{N},\rho_{n}]}\right\rVert_{\operatorname{Tr}}=\mathcal{O}\left(\frac{\lVert\hat{N}\rVert_{\infty}^{2}}{M}\right)\ ,$ (S149)
from Eq. (S25). The flow duration is chosen to be $s_{n}=\mathcal{O}(\lVert\hat{D}\rVert^{-1})$ as in Lemma 23, which makes the error scaling independent of the rescaling of $\hat{D}$. Therefore, one can make the locally accurate implementation, by increasing the number of copies $M$.
@@
V.4 Quantum imaginary-time evolution
Up to this point, we have focused on the diagonalization of the state $\rho_{n}$ (or $\hat{P}_{n}$) as a result of double-bracket iterations. However, Lemma 23 tells us an additional property: the fixed-point is not only diagonal in the basis of $\hat{D}$, but its eigenvalues are also rearranged in an non-increasing order. In other words, the eigenvector of $\rho_{0}$ with the highest eigenvalue will align with the eigenvector of $\hat{D}$ that has the lowest eigenvalue.
Now suppose that $\rho_{0}=|{\psi}\rangle\!\langle{\psi}|$ is a pure state. Then the highest eigenvalue of it is $1$ with the corresponding eigenvector $|{\psi}\rangle$. Then, as the result of the double-bracket iterations, $\rho_{\infty}$ will become an eigenvector of $\hat{D}$ with the largest eigenvalue. If we set $\hat{D}=-\hat{H}$, this means that the double-bracket iteration prepares the ground state of the Hamiltonian $\hat{H}$.
In Ref. [6], it has been found that this process is equivalent to the discretized version of quantum imaginary-time evolution. The goal is to apply an operator similar to the usual Hamiltonian evolution $e^{-it\hat{H}}$ but with an imaginary-time $t=-i\beta$. Clearly, this is not a unitary operator, and will not result in a quantum state after the application. Hence, we instead need to consider a unitary flow of a pure state $|{\psi(\beta)}\rangle\propto e^{-\beta H}|{\psi(0)}\rangle$ starting $\beta=0$ and evolving towards $\beta\to\infty$. The corresponding differential equation is then written as
$\partial_{\beta}|{\psi(\beta)}\rangle=-(\hat{H}-E_{\beta})|{\psi(\beta)}\rangle\propto-\hat{H}|{\psi(\beta)}\rangle\ ,$ (S150)
where $E_{\beta}=\langle{\psi(\beta)}|\hat{H}|{\psi(\beta)}\rangle$. Surprisingly, this equation is equivalent to another equation
$\partial_{\beta}|{\psi(\beta)}\rangle=\left[|{\psi(\beta)}\rangle\!\langle{\psi(\beta)}|,\hat{H}\right]|{\psi(\beta)}\rangle\ ,$ (S151)
which can be discretized into finite steps with size $s_{n}$ as
$|{\psi_{n+1}}\rangle=e^{s_{n}[|{\psi_{n}}\rangle\!\langle{\psi_{n}}|,\hat{H}]}|{\psi_{n}}\rangle\ .$ (S152)
This comes from the fact that the cost function $f$ in Eq. (S141) can be written as $f(|{\psi}\rangle\!\langle{\psi}|)=1+\operatorname{Tr}[\hat{H}^{2}]+2\langle{\psi}|\hat{H}|{\psi}\rangle$, and the double-bracket iteration, which is a gradient flow in terms of this function, is effectively giving the gradient flow in terms of the average energy of the pure state $|{\psi}\rangle$. Note that Eq. (S152) is exactly in the form of Eq. (S140) with $\hat{D}=-\hat{H}$ and $\hat{P}_{n}=|{\psi_{n}}\rangle\!\langle{\psi_{n}}|$.
Furthermore, Theorem 2 of Ref. [6] guarantees the fast spectral convergence of this algorithm. Therefore, we can apply Thm. 12, which ensures the efficient QDP implementation.
One potential difficulty of implementing the memory-call $e^{s_{n}[|{\psi_{n}}\rangle\!\langle{\psi_{n}}|,\hat{H}]}$ directly from memory-usage queries is the requirement for a prior information about the Hamiltonian $\hat{H}$. Hence, in the main text, we considered a sampling based Hamiltonian simulation scenario studied in Refs. [13, 16]. In such settings, many copies of the state $\chi\propto\hat{H}-\lambda_{0}\mathmybb{1}$ are given. Then, instead of considering the single variable function $\hat{\mathcal{N}}(\rho)=-i[\rho,\hat{H}]$ as the one defining the memory-call, we can work with a two-variable function $f(\rho,\chi)=-i[\rho,\chi]$. Such memory-usage queries can be made using the ideas from Lemma 4.
V.5 Oblivious Schmidt decomposition
@@
$|{\psi}\rangle_{AB}=\sum_{j}\sqrt{\lambda_{j}}|{\phi_{j}}\rangle_{A}|{\chi_{j}}\rangle_{B}\ ,$ (S153)
@@
$e^{i\mathcal{N}(\psi_{n})}=e^{s_{n}[\hat{D},\rho_{n}^{(A)}]}\otimes\mathmybb{1}_{B}\ ,$ (S154)
@@
^$\hat{\mathcal{E}}^{[\mathcal{N},\psi_{A_{1}B_{1}}]}_{r}(\sigma_{A_{2}B_{2}})=\operatorname{Tr}_{A_{1}B_{1}}\left[e^{-ir\hat{N}}(\psi_{A_{1}B_{1}}\otimes\sigma_{A_{2}B_{2}})e^{ir\hat{N}}\right]\ ,$ (S155)
where $\hat{N}=is_{n}\sum_{jk}(\mu_{k}-\mu_{j})|{kj}\rangle\!\langle{jk}|_{A_{1}A_{2}}\otimes\mathmybb{1}_{B_{1}B_{2}}$ and $\{\mu_{j}\}_{j}$ are the diagonal entries of $\hat{D}$. Each query ^$\hat{\mathcal{E}}^{[\mathcal{N},\psi_{A_{1}B_{1}}]}_{r}(\sigma_{A_{2}B_{2}})$ consumes a copy of $\psi_{A_{1}B_{1}}$. Furthermore, following the same calculation from the previous subsection, we obtain
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{[\mathcal{N},\psi_{A_{1}B_{1}}]}_{1/M}\right)^{M}-e^{i\mathcal{N}(\psi_{n})}\right\rVert_{\operatorname{Tr}}=\mathcal{O}\left(\frac{\lVert\hat{N}\rVert_{\infty}^{2}}{M}\right)\ ,$ (S156)
@@
If recursions are implemented with sufficiently good precision, Theorem 1 in the main text holds and we approach the fixed-point of the recursion, which is a diagonal matrix $\rho_{\infty}^{(A)}$ isospectral to the initial reduced state $\rho_{0}^{(A)}$. In other words, $\rho_{N}^{(A)}$ with some large number $N$ becomes close to the desired diagonal matrix
$\rho_{N}\simeq\sum_{j}\lambda_{j}|{j}\rangle\!\langle{j}|_{A}\ ,$ (S157)